OpenStudy (anonymous):

Simplify: (2y)^3 * y^-1 can someone help me?

OpenStudy (solomonzelman):

\((2y)^3\times y^{-1}\)

OpenStudy (solomonzelman):

at first, \(2^3\times y^3\times y^{-1}\)

OpenStudy (solomonzelman):

then \(2^3\times y^{3+ (-1)}\)

OpenStudy (solomonzelman):

please finish it from here